area of rhombus is 10 sq. unit. its diagonal intersect at (0,0).if one vertex is (3,4) then one of the other vertices can be: pair of straight line question

Grade:11

1 Answers

Rajshree
6 years ago
Area=1/2d1×d220=d1×d2Let the rhombus ABCDN the intersection point is DPoint be C(3,4)CD=5 #applying distance D1=10D2=2Now. Its easy u can